As indicated in the chapter, return on investment (ROI) is well entrenched in business practice. However, its use can have negative incentive effects on managerial behavior. For example, assume you are the manager of an investment center and that your annual bonus is a function of achieved ROI for your division. You have the opportunity to invest in a project that would cost $420,000 and that would increase annual operating income of your division by $30,000. (This level of return is considered acceptable from top management’s standpoint.) Currently, your division generates annual operating profits of approximately $705,000, on an asset base (i.e., level of investment) of $4,500,000.

Required:
1.

What is the current return on investment (ROI) being realized by your division (i.e., before considering the new investment)? (Round your answer to 2 decimal places.)

      

2.

What would happen to the near-term ROI of your division after adding the effect of the new investment? (Round your answer to 2 decimal places. (i.e. .1234 = 12.34%))

      

3.

As manager of this division, given your incentive-compensation plan, would you be motivated to make the new investment?

  • Yes

  • No
Answer #1
1] ROI = 705000/4500000 = 15.67%
2] ROI = (705000+30000)/(4500000+420000) = 14.94%
3] NO.
Reason is that if the new project is undertaken the
overall ROI of the division will go down affecting the
amount of annual bonus adversely.
Answer #2

To calculate the answers, we need to perform the following calculations:

  1. Current Return on Investment (ROI): ROI = (Operating Income / Investment) * 100 ROI = ($705,000 / $4,500,000) * 100 ROI ≈ 15.67%

  2. Near-term ROI after adding the effect of the new investment: New Operating Income = Current Operating Income + Increase in Operating Income New Operating Income = $705,000 + $30,000 = $735,000

    New Investment = Current Investment + Cost of New Investment New Investment = $4,500,000 + $420,000 = $4,920,000

    New ROI = (New Operating Income / New Investment) * 100 New ROI = ($735,000 / $4,920,000) * 100 New ROI ≈ 14.92%

  3. As the manager, let's compare the current ROI with the near-term ROI after the new investment: Current ROI = 15.67% New ROI = 14.92%

    Since the new investment would result in a lower ROI, and your annual bonus is a function of achieved ROI, you would not be motivated to make the new investment. Therefore, the answer is "No."


answered by: mervetokaz